OpenStudy (anonymous):

Find the x: 2x^2+x=3x

OpenStudy (turingtest):

step 1) get all the x's together on one side of the equals sign

OpenStudy (anonymous):

can you just do the whole math? please? :(

OpenStudy (turingtest):

(this is called collecting like terms) -no, sorry I prefer teaching to pumping out answers

OpenStudy (anonymous):

i dont have time look PLEASE!!

OpenStudy (anonymous):

2x^2-2x=0 2x(x-1)=0 x=0,1

OpenStudy (anonymous):

Solve the X here too :( I soo hate maths!! :'( x^2-12x-64=0

OpenStudy (anonymous):

I am a newbie here but I wish there was an archive of sorts here for all to read and learn from the great answers posted previously ( I have seen some of yours @TuringTest ). There are quite a few guys patiently answering here, but a whole lot more looking for just quick solutions. I hope in all this everyone is learning something at their speed.